# Accelerating Goal-Conditioned Reinforcement Learning Algorithms and Research

### v3

@@ B.1 Environment details
Pusher Easy 0.1 No Generalized x coordinate sampled from $[-0.55,-0.25]$, y coordinate sampled from $[-0.2,0.2]$
Pusher Hard 0.1 No Generalized x coordinate sampled from $[-0.65,0.35]$, y coordinate sampled from $[-0.55,0.45]$
@@ B.2 Benchmark details
Our experiments use JaxGCRL suite of simulated environments described in Section 4.2. We evaluate algorithms in an online setting, with a UTD ratio $1$:$16$ for CRL, TD3, TD3+HER, SAC, SAC+HER, and $1:5$ for PPO. We use a batch size of $256$ and a discount factor of $0.99$ for all methods except PPO, for which we use a discount factor of $0.97$. For every environment, we sample evaluation goals from the same distribution as training ones and use a replay buffer of size $10$M for CRL, TD3, TD3+HER, SAC, and SAC+HER. We use $1024$ parallel environments for all methods except for PPO, where we use $4096$ parallel environments to collect data. All experiments are conducted for $50$ million environment steps.
@@ B.3 Benchmark parameters
The parameters used for benchmarking experiments can be found in Table 2.
@@ Appendix C Random Goals
Our loss in Eq. 3 differs from the original CRL algorithm (Eysenbach et al.,, 2022) by sampling goals from the same trajectories as states during policy extraction, rather than random goals from the replay buffer. Mathematically, we can generalize Eq. 3 to account for either of these strategies by adding a hyperparameter $\alpha$ to the loss controlling the degree of random goal sampling during training.
@@
+\,+ $\alpha\cdot\mathbb{E}_{p(s,a)p({\color[rgb]{0.75,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0,0}g})\pi_{\theta}(a^{\prime}|s,{\color[rgb]{0.75,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0,0}g})}\left[f_{\phi,\psi}(s,a^{\prime},{\color[rgb]{0.75,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0,0}g})\right]$
The hyperparameter $\alpha$ controls the rate of counterfactual goal learning, where the policy is updated based on the critic’s evaluation of goals that did not actually occur in the trajectory. We find that taking, $\alpha=0$ (i.e., no random goal sampling) leads to better performance, and suggest using the policy loss in Eq. 3 for training contrastive RL methods.
@@ D.1 Q-function is probability
This proof follows closely the one presented in Eysenbach et al., (2022).
@@
On the second line, we have used the Markov property. We can now substitute this into Eq. 20:
@@ Appendix E Failed Experiments
    Weight Decay: Prior work (Nauman et al., 2024b, ) indicated that regularizing critic weights might improve learning stability. We did not observe significant upgrades in the CRL setup, perhaps due to a much lower ratio of updates per environment step. We tested this only for small architectures.
@@
    Random Goals: In prior implementation (Eysenbach et al.,, 2022) using random goals in actor loss resulted in higher performance. We did not observe that in our online setting.
@@ Appendix F Pseudocode
Pseudocode for the contrastive learning algorithms studied is presented in Algorithm 1.
@@
[…]
3:Observe state $s$ and sample an action $a\sim\pi(s,g)$
4:Execute $a$ in the environment
5:Observe next state $s^{\prime}$ and done signal $d$ to indicate whether $s^{\prime}$ is terminal
[…]
7:Reset environment state and sample new goal
8:Store current trajectory for this environment in $\mathcal{D}$
[…]
10: Randomly sample (with discount) a batch $\mathcal{B}$ from $\mathcal{D}$ of state-action pairs and goals from their future
11: Update critic: $(\phi,\psi)\leftarrow(\phi,\psi)-\alpha\nabla_{\phi,\psi}\bigl{[}\mathcal{L}_{\text{Critic}}(\mathcal{B};\phi,\psi)+\beta\mathcal{L}_{\text{logsumexp}}(\mathcal{B},\phi,\psi)\bigr{]}$
[…]

### v4

@@ B.1 Environment details
Pusher Easy 0.1 No Generalized x coordinate sampled from [0.55,0.25][-0.55,-0.25], y coordinate sampled from [0.2,0.2][-0.2,0.2]
Pusher Hard 0.1 No Generalized x coordinate sampled from [0.65,0.35][-0.65,0.35], y coordinate sampled from [0.55,0.45][-0.55,0.45]
@@ B.2 Benchmark details
Our experiments use JaxGCRL suite of simulated environments described in Section˜4.2. We evaluate algorithms in an online setting, with a UTD ratio 11:$16$ for CRL, TD3, TD3+HER, SAC, SAC+HER, and $1:5$ for PPO. We use a batch size of $256$ and a discount factor of $0.99$ for all methods except PPO, for which we use a discount factor of $0.97$. For every environment, we sample evaluation goals from the same distribution as training ones and use a replay buffer of size 1010M for CRL, TD3, TD3+HER, SAC, and SAC+HER. We use $1024$ parallel environments for all methods except for PPO, where we use $4096$ parallel environments to collect data. All experiments are conducted for $50$ million environment steps.
@@ B.3 Benchmark parameters
The parameters used for benchmarking experiments can be found in Table˜2.
@@ Appendix C Random Goals
Our loss in Eq.˜3 differs from the original CRL algorithm (eysenbach2022contrastive) by sampling goals from the same trajectories as states during policy extraction, rather than random goals from the replay buffer. Mathematically, we can generalize Eq.˜3 to account for either of these strategies by adding a hyperparameter $\alpha$ to the loss controlling the degree of random goal sampling during training.
@@
$+\,$ $\alpha\cdot\mathbb{E}_{p(s,a)p({\color[rgb]{0.75,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0,0}g})\pi_{\theta}(a^{\prime}|s,{\color[rgb]{0.75,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0,0}g})}\left[f_{\phi,\psi}(s,a^{\prime},{\color[rgb]{0.75,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0,0}g})\right]$
The hyperparameter $\alpha$ controls the rate of counterfactual goal learning, where the policy is updated based on the critic’s evaluation of goals that did not actually occur in the trajectory. We find that taking, $\alpha=0$ (i.e., no random goal sampling) leads to better performance, and suggest using the policy loss in Eq.˜3 for training contrastive RL methods.
@@ D.1 Q-function is probability
This proof follows closely the one presented in eysenbach2022contrastive.
@@
On the second line, we have used the Markov property. We can now substitute this into Eq.˜20:
@@ Appendix E Failed Experiments
    Weight Decay: Prior work (nauman2024bigger) indicated that regularizing critic weights might improve learning stability. We did not observe significant upgrades in the CRL setup, perhaps due to a much lower ratio of updates per environment step. We tested this only for small architectures.
@@
    Random Goals: In prior implementation (eysenbach2022contrastive) using random goals in actor loss resulted in higher performance. We did not observe that in our online setting.
@@ Appendix F Pseudocode
Pseudocode for the contrastive learning algorithms studied is presented in Algorithm˜1.
@@
2: Initialize $\phi$, $\psi$, $\pi$ and an empty replay buffer $\mathcal{D}$
3: repeat
4:   in parallel over environments
5:    Observe state $s$ and sample an action $a\sim\pi(s,g)$
6:    Execute $a$ in the environment
7:    Observe next state $s^{\prime}$ and done signal $d$ to indicate whether $s^{\prime}$ is terminal
8:    Append $(s,a,s^{\prime})$ to current trajectory for this environment
9:    if $s^{\prime}$ is terminal then
10:      Reset environment state and sample new goal
11:      Store current trajectory for this environment in $\mathcal{D}$
12:      Start new trajectory for this environment
13:   for $j=1,\ldots,\texttt{num\_updates}$ do
14:    Randomly sample (with discount) a batch $\mathcal{B}$ from $\mathcal{D}$ of state-action pairs and goals from their future
15:    Update critic: $(\phi,\psi)\leftarrow(\phi,\psi)-\alpha\nabla_{\phi,\psi}\bigl[\mathcal{L}_{\text{Critic}}(\mathcal{B};\phi,\psi)+\beta\mathcal{L}_{\text{logsumexp}}(\mathcal{B},\phi,\psi)\bigr]$
16:    Update policy: $\pi\leftarrow\pi-\alpha\nabla_{\pi}\bigl[\mathcal{L}_{\text{Actor}}(\mathcal{B};\phi,\psi,\pi)\bigr]$
17: until convergence